# ExLM: Rethinking the Impact of [MASK] Tokens in Masked Language Models

### v3

@@ 4.1 Overview
Through the analysis in Section 3, we have identified that corrupted semantics is the main factor affecting MLM’s performance. When semantics of the context are severely corrupted, it leads to a significant loss of information in the context, making it increasingly difficult for the MLM to restore the original tokens.
Building on this, an intuitive approach to improving MLM is: how can we enhance the model’s ability to predict the right and complete semantic information from incomplete contexts? The missing semantic information can be summarized into two parts: one is the semantics of the missing tokens themselves, and the other is the semantic dependencies between these missing tokens. Following this core idea, we propose a novel enhanced-context MLM (ExLM), with two main improvements:
  • States Expansion: For each [MASK] token in the input context, multiple hidden states are created, enlarging the semantic space for the model, thereby capturing richer semantic information and improving the model’s ability to predict missing semantic information.
  • Dependency Capture: A transition matrix is used to explicitly model the semantic dependencies between the states, and a States Alignment algorithm based on dynamic programming efficiently provides supervision signals to these hidden states.
In this way, by predicting richer semantic information and explicitly modeling the semantic dependencies between the missing information, the ExLM gains enhanced modeling capabilities, allowing it to better restore the missing semantic information from contexts with corrupted semantics. Therefore, in this section, we will first introduce the model details for context enhancement (Section 4.2) and then present the pre-training algorithm for ExLM, namely States Alignment (Section 4.3).
@@ 4.3 Pre-training of ExLM: States Alignment
A key challenge in training ExLM lies in the fact that each [MASK] token is expanded into multiple hidden states. This expansion means there are more hidden states than there are target tokens to predict. Consequently, we must determine an alignment between these states and the target tokens—that is, deciding which hidden state should be responsible for predicting which token. This alignment process is at the heart of our States Alignment algorithm.
Formulating States Alignment as a DAG decoding problem.
This objective requires summing over all possible alignment paths $\mathbf{A}\in\Gamma$, which can be computationally expensive. To address this, we employ a dynamic programming (DP) algorithm that efficiently computes this sum with a complexity of $\mathcal{O}(M)$, where $M$ is the number of all masked tokens. This approach significantly improves computational efficiency. Details of the DP algorithm are provided in Appendix H and the efficiency analysis is provided in Appendix I.
5 Experimental Results
We pre-train ExLM on data of text and SMILES separately. Then we fine-tune and evaluate ExLM across diverse benchmarks and verify the contribution of each component through ablation studies. Finally, a visualization analysis is included to explain the advantages of unified modeling.
@@ 5 Experimental Results
5.1 Results on SMILES Representation Learning
Pre-training.
For SMILES pre-training, we use the large-scale molecular dataset provided by Zhou et al. (2023), which includes SMILES information for $19$ million molecules. We tokenize SMILES sequences with the regular expression from Schwaller et al. (2018). The pre-training hyperparameters can be found in Appendix J.
Fine-tuning.
For fine-tuning, we employ the widely-recognized MoleculeNet benchmark (Wu et al., 2018). We follow the same data split as used by Zhou et al. (2023). Details of the fine-tuning datasets and baselines can be found in Appendix L. We fine-tune the ExLM model on downstream task datasets using three different random seeds and reported the average performance of the model.
Results.
5.2 Results on Textual Representation Learning
Pre-training.
For textual pre-training, we adopt the English Wikipedia and BookCorpus datasets (Devlin, 2018) as the pre-training dataset. The model size of ExLM is consistent with BERT base (Devlin, 2018). We also train ExLMLARGE with the same size as BERT large. For more details about pre-training settings, please see Appendix N.
Fine-tuning.
We evaluate the ExLM model using the the GLUE (Wang et al., 2018) and SQuAD 2.0 (Rajpurkar et al., 2018) benchmarks. Detailed of the GLUE benchmark are provided in Appendix M. For fine-tuning, we follow the standard procedures used in BERT (Devlin, 2018) and RoBERTa (Liu, 2019). We also provide the hyperparameter search space for fine-tuning in Appendix O. Consistent with previous studies (Liu, 2019), all reported fine-tuning results represent the median values obtained from five different random seeds across both GLUE and SQuAD benchmarks.
Results.
[…]
ExLM w/o 2D RoPE 84.6 91.1 91.3 56.7 80.9
ExLM w/o Transitions 83.8 90.9 91.1 55.6 80.4
Vanilla MLM 83.6 90.0 90.3 54.7 79.6
Vanilla MLM++ 84.4 91.2 90.6 56.3 80.7
ExLM 85.1 91.4 91.3 57.6 81.4
@@ 5.3 Analytical Experiments of ExLM
We validate ExLM’s effectiveness through ablation studies and explore the impact of expanded states and mask ratios on its performance. A case study also demonstrates its good ability to capture rich semantic information. Efficiency and entropy analysis of ExLM are provided in Appendix Q.
@@
We conduct a case study to evaluate the model’s performance on specific inputs, as shown in Figure 7. The results demonstrate that the model effectively handles multimodal semantic information, capturing four possible semantic choices for each [MASK] token. The transition matrix models the relationships between these choices, for example, when the first [MASK] token is amazing, the second is more likely to be glad, and when the first is terrible, the second tends to be sorry. A more detailed analysis can be found in Appendix P. These results show that ExLM can capture rich contextual information.
@@ 6 Conclusions
Acknowledgements
We would like to thank Jie Zhu from University of Oxford and Liang Ding from the JD Explore Academy for their insightful discussions on the project. We also thank all other members from Dlib in PKU for their valuable feedback given during the internal discussions. This paper is partially supported by National Natural Science Foundation of China with Grant Number 62276002.
@@ Appendix A Related Works
Currently, language models serve as an important role in the self-supervised representation learning field and achieve excellent performance across a wide range of tasks. The ELMo model (Peters et al., 2018) first introduces a self-supervised language representation learning model based on bidirectional LSTM (Hochreiter, 1997). Later, the pre-trained masked language model based on the transformer architecture, BERT (Devlin, 2018), is proposed in the field of text modeling. BERT uses a bidirectional transformer encoder to extract meaningful semantic representations from large amounts of text data and applies them to various downstream tasks such as text understanding (Devlin, 2018) and text correction (Zhang et al., 2020; Dai et al., 2022). With the help of the transformer architecture (Vaswani et al., 2017) and the MLM training approach, this model achieves great success in the text domain. Following this, a large number of follow-up works aim to improve these bidirectional encoder-based masked language models (Liu, 2019; Lan, 2019; Ke et al., 2020; Joshi et al., 2020; He et al., 2020; Shin et al., 2020; Du et al., 2021; Fu et al., 2022; Wettig et al., 2022; Meng et al., 2022, 2023; He et al., 2022b). Moreover, MLM also succeeds in various representation learning tasks, including images (Bao et al., 2021; Xie et al., 2022; He et al., 2022a), videos (Tong et al., 2022; Wang et al., 2022), codes (Feng et al., 2020; Guo et al., 2020; He et al., 2021), small molecules(Wang et al., 2019; Chithrananda et al., 2020; Ross et al., 2022; Pan, 2023; Feng et al., 2024; Yang et al., 2024; Zheng et al., 2024a), proteins (Elnaggar et al., 2021; Lin et al., 2022, 2023; Zheng et al., 2024b; Hayes et al., 2025), and single-cell data (Yang et al., 2022; Theodoris et al., 2023).
@@ Appendix C Expectation and Variance of the Proportion of Corrupted Semantics
This analysis leverages the independence of masking each token, leading to the conclusion that since the length of the input sequence $N$ is typically large, the variance of $s$ is very small.
@@ Appendix D A Brief Introduction to SMILES
SMILES (Simplified Molecular Input Line Entry System) (Weininger, 1988) is a notation system used to represent chemical structures in a text format. It encodes molecular information using a series of characters, where atoms are represented by their chemical symbols (e.g., C for carbon, O for oxygen), bonds by symbols like ”-” (single bond), ”=” (double bond), and ”#” (triple bond), and branches are indicated with parentheses. For example, the SMILES representation of ethanol is ”CCO”, where ”C” stands for carbon atoms and ”O” represents an oxygen atom, with a single bond between them. The notation captures the connectivity between atoms, bond types, and sometimes stereochemistry, making it a compact and easily readable way to represent molecular structures. The principle behind SMILES is that it provides a linear representation of the molecule that can be easily interpreted by both humans and computational systems, facilitating tasks like database searching, molecular simulations, and chemical informatics.
@@ Appendix G Training Curves in the Repeated MLM Experiments
Appendix H Dynamic Programming for Efficient Alignment
Naively summing over all paths $\mathbf{A}\in\Gamma$ can be prohibitively expensive, since the number of possible paths grows exponentially. To solve this, we use a dynamic programming (DP) approach, which is also adopted by the training algorithms for connectionist temporal classification (Graves et al., 2006; Saharia et al., 2020) and DA-Transformer (Huang et al., 2022; Shao et al., 2022; Huang et al., 2023, 2024).
In this DP scheme, we define $f_{i,u}$ as the cumulative probability of all partial paths ending at state $u$ (a node in the DAG) that have generated the first $i$ tokens of $\mathbf{Y}$. Formally, $u$ indexes the states in our DAG in a manner that respects the acyclic property (i.e., we only move forward in the state sequence), and $i$ ranges over the positions of the target sequence $\mathbf{Y}$.
By repeatedly computing $f_{i,u}$ for all $i$ and $u$, we ultimately obtain $f_{M,L}$ (where $M$ is the length of the target sequence $\mathbf{Y}$ (equal to the number of masked tokens), and $L$ indexes the final state.). The final training objective is:
$\mathcal{L}_{\text{SA}}\;=\;-\,\log f_{M,\,L}.$
This dynamic programming solution reduces the alignment problem’s time complexity to $\mathcal{O}(M\times L^{2})$, which is considerably more efficient than enumerating all possible paths. Moreover, it can be implemented with a manageable number of high-level operations (e.g., $\mathcal{O}(M)$ PyTorch operations), making it practical for most training scenarios.
Appendix I CUDA-Accelerated Dynamic Programming Framework: Details and Efficiency Analysis
We adopted an algorithm provided by DA-Transformer(Huang et al., 2022), utilizing CUDA-accelerated dynamic programming to efficiently perform the states alignment pre-training. The core of the algorithm utilizes CUDA (Compute Unified Device Architecture) to parallelize the computations involved in dynamic programming for sequence modeling. The primary objective is to efficiently calculate forward ($\alpha$) and backward ($\beta$) probabilities, which are crucial for evaluating sequence likelihoods and performing gradient-based optimization during model training. By leveraging the parallel processing capabilities of GPUs, the algorithm distributes computations across numerous threads and blocks, significantly reducing runtime compared to traditional CPU-based implementations.
I.1 Dynamic Programming Framework: Forward Probability ($\alpha$) and Backward Probability ($\beta$)
@@ I.2 Gradient Calculations
I.3 Time Complexity Analysis
@@
I.4 Space Complexity Analysis
@@
Appendix J Hyper-Parameter Configuration for Molecular Pre-training
@@
Appendix K Hyper-Parameter Configuration for Molecular Fine-tuning
@@
Appendix L Details of Molecular Fine-tuning Datasets and Baselines
@@
Appendix M Details of the GLUE Benchmark
Below are detailed descriptions of all the GLUE tasks:
@@
Appendix N Hyper-Parameter Configuration for Textual Pre-training
@@
Appendix O Hyper-Parameter Configuration for Textual Fine-tuning
@@
Appendix P Detailed Case Study
@@
Appendix Q More Analytical Experiments of ExLM
Q.1 Entropy Analysis on ExLM
@@
Q.2 Training Efficiency of ExLM
@@
Q.3 Performance of Sparse ExLM

### v5

@@ 4.1 Overview
Through the analysis in Section 3, we have identified that corrupted semantics is the main factor affecting MLM’s performance. When semantics of the context are severely corrupted, the resulting multimodality makes it increasingly difficult for the MLM to restore the original tokens.
Building on this, an intuitive approach to improving the MLM is: how can we enhance the model’s ability to better handle semantic multimodality? More specifically, the semantic multimodality can be divided into two aspects:
  • Intra-token Multimodality: The potential choices for each missing token become more diverse, and the significant semantic differences among these choices increase the semantic diversity and ambiguity.
  • Inter-token Multimodality: The meaning of one token is intricately linked to and influenced by the meanings of other tokens, resulting in complex semantic interactions and dependencies between missing tokens.
Therefore, we need to enhance the model’s capabilities in two areas: one is improving its ability to model diverse and ambiguous semantics, and the other is enhancing its ability to capture semantic dependencies between missing tokens. Following this core idea, we propose a novel enhanced-context MLM (ExLM) with two main improvements:
  • States Expansion: For each [MASK] token in the input context, multiple hidden states are created, enlarging the semantic space for the model. A larger semantic space enables the model to capture richer and more diverse semantic information, better handling the semantic diversity introduced by multimodality.
  • Dependency Capture: A transition matrix is used to explicitly model the semantic dependencies between different states, and the States Alignment algorithm based on dynamic programming efficiently provides supervision signals to these hidden states.
Therefore, we will first introduce the details of context enhancement, which includes States Expansion and Dependency Capture (Section 4.2), and then present the States Alignment algorithm for ExLM pre-training (Section 4.3).
@@ 4.3 Pre-training of ExLM: States Alignment
A key challenge in training ExLM lies in the fact that each [MASK] token is expanded into multiple hidden states. This expansion means that there are more hidden states than the target tokens to predict. Consequently, we must determine an alignment between these states and the target tokens—that is, deciding which hidden state should be responsible for predicting which token. This alignment process is at the heart of our States Alignment algorithm.
@@
Formulating States Alignment as a DAG decoding problem.
This objective requires summing over all possible alignment paths $\mathbf{A}\in\Gamma$, which can be computationally expensive. To address this, we employ a dynamic programming (DP) algorithm that efficiently computes this sum with a complexity of $\mathcal{O}(M)$, where $M$ is the number of all masked tokens.
More specifically, we adopted a dynamic programming algorithm similar to that used in DA-Transformer (Huang et al., 2022). In this DP scheme, we define $f_{i,u}$ as the cumulative probability of all partial paths ending at state $u$ (a node in the DAG) that have generated the first $i$ tokens of $\mathbf{Y}$. Formally, $u$ indexes the states in our DAG in a manner that respects the acyclic property (i.e., we only move forward in the state sequence), and $i$ ranges over the positions of the target sequence $\mathbf{Y}$.
By computing $f_{i,u}$ across all $i$ and $u$, we eventually obtain $f_{M,L}$, where $M$ is the number of masked tokens in the target sequence $\mathbf{Y}$, and $L$ refers to the final state index. The final training objective is:
$\mathcal{L}_{\text{SA}}\;=\;-\,\log f_{M,\,L}.$
This dynamic programming approach reduces the time complexity of the alignment problem to $\mathcal{O}(M\times L^{2})$, offering a significant improvement over exhaustive path enumeration. In practice, the computation can be further optimized to $\mathcal{O}(M)$ by leveraging parallelized operations provided by PyTorch (Paszke, 2019), making the method highly efficient and suitable for large-scale training. A detailed analysis of the DP algorithm’s efficiency is provided in Appendix H.
5 Experimental Results
We pre-train ExLM [^1] Code is released at https://github.com/zhengkangjie/ExLM. on data of text and SMILES separately. Then we fine-tune and evaluate ExLM across diverse benchmarks and verify the contribution of each component through ablation studies. Finally, a visualization analysis is included to explain the advantages of unified modeling.
@@ 5 Experimental Results
5.1 Results on SMILES Representation Learning
Pre-training.
For SMILES pre-training, we use the large-scale molecular dataset provided by Zhou et al. (2023), which includes SMILES information for $19$ million molecules. We tokenize SMILES sequences with the regular expression from Schwaller et al. (2018). The pre-training hyperparameters can be found in Appendix I.
Fine-tuning.
For fine-tuning, we employ the widely-recognized MoleculeNet benchmark (Wu et al., 2018). We follow the same data split as used by Zhou et al. (2023). Details of the fine-tuning datasets and baselines can be found in Appendix K. We fine-tune the ExLM model on downstream task datasets using three different random seeds and reported the average performance of the model.
Results.
5.2 Results on Textual Representation Learning
Pre-training.
For textual pre-training, we adopt the English Wikipedia and BookCorpus datasets (Devlin, 2018) as the pre-training dataset. The model size of ExLM is consistent with BERT base (Devlin, 2018). We also train ExLMLARGE with the same size as BERT large. For more details about pre-training settings, please see Appendix M.
Fine-tuning.
We evaluate the ExLM model using the the GLUE (Wang et al., 2018) and SQuAD 2.0 (Rajpurkar et al., 2018) benchmarks. Detailed of the GLUE benchmark are provided in Appendix L. For fine-tuning, we follow the standard procedures used in BERT (Devlin, 2018) and RoBERTa (Liu, 2019). We also provide the hyperparameter search space for fine-tuning in Appendix N. Consistent with previous studies (Liu, 2019), all reported fine-tuning results represent the median values obtained from five different random seeds across both GLUE and SQuAD benchmarks.
Results.
[…]
ExLM w/o 2D RoPE 84.6 91.1 91.3 56.7 80.9
ExLM w/o Transitions 83.8 90.9 91.1 55.6 80.4
Vanilla MLM 83.6 90.0 90.3 54.7 79.6
Vanilla MLM++ 84.4 91.2 90.6 56.3 80.7
ExLM 85.1 91.4 91.3 57.6 81.4
@@ 5.3 Analytical Experiments of ExLM
We validate ExLM’s effectiveness through ablation studies and explore the impact of expanded states and mask ratios on its performance. A case study also demonstrates its good ability to capture rich semantic information. Efficiency and entropy analysis of ExLM are provided in Appendix P.
Ablation studies on ExLM.
@@
We conduct a case study to evaluate the model’s performance on specific inputs, as shown in Figure 7. The results demonstrate that the model effectively handles multimodal semantic information, capturing four possible semantic choices for each [MASK] token. The transition matrix models the relationships between these choices, for example, when the first [MASK] token is amazing, the second is more likely to be glad, and when the first is terrible, the second tends to be sorry. A more detailed analysis can be found in Appendix O. These results show that ExLM can capture rich contextual information.
@@ 6 Conclusions
Acknowledgements
This paper is partially supported by grants from the National Key Research and Development Program of China with Grant No. 2023YFC3341203 and the National Natural Science Foundation of China (NSFC Grant Number 62276002). We would like to thank Jie Zhu from University of Oxford and Liang Ding from the JD Explore Academy for their insightful discussions on the project. We also thank all other members from Dlib in PKU for their valuable feedback given during the internal discussions.
@@ Appendix A Related Works
Currently, language models serve as an important role in the self-supervised representation learning field and achieve excellent performance across a wide range of tasks. The ELMo model (Peters et al., 2018) first introduces a self-supervised language representation learning model based on bidirectional LSTM (Hochreiter, 1997). Later, the pre-trained masked language model based on the transformer architecture, BERT (Devlin, 2018), is proposed in the field of text modeling. BERT uses a bidirectional transformer encoder to extract meaningful semantic representations from large amounts of text data and applies them to various downstream tasks such as text understanding (Devlin, 2018) and text correction (Zhang et al., 2020; Dai et al., 2022; Zheng et al., 2023). With the help of the transformer architecture (Vaswani et al., 2017) and the MLM training approach, this model achieves great success in the text domain. Following this, a large number of follow-up works aim to improve these bidirectional encoder-based masked language models (Liu, 2019; Ke et al., 2020; Shin et al., 2020; Namazifar et al., 2021; Du et al., 2021; Fu et al., 2022; Wettig et al., 2022; Meng et al., 2022, 2023; He et al., 2022b). Moreover, MLM also succeeds in various representation learning tasks, including images (Bao et al., 2021; Xie et al., 2022; He et al., 2022a), videos (Tong et al., 2022; Wang et al., 2022), codes (Feng et al., 2020; Guo et al., 2020; He et al., 2021), small molecules(Wang et al., 2019; Chithrananda et al., 2020; Ross et al., 2022; Pan, 2023; Feng et al., 2024; Yang et al., 2024; Zheng et al., 2024a), proteins (Elnaggar et al., 2021; Lin et al., 2022, 2023; Zheng et al., 2024b; Hayes et al., 2025), and single-cell data (Yang et al., 2022; Theodoris et al., 2023).
@@ Appendix C Expectation and Variance of the Proportion of Corrupted Semantics
This analysis leverages the independence of masking each token. It proves that when the repetition times are $k$ and the mask ratio is $p$, the proportion of corrupted semantics is $p^{k}$. Moreover, because the input sequence length $N$ is typically large, the variance of $s$ becomes vanishingly small. Concretely, let’s take:
$p=0.387,\quad k=2,\quad N=512.$
Then
$\mathbb{E}[s]=p^{k}=0.387^{2}\approx 0.1498,$
$\mathrm{Var}(s)=\frac{0.1498\,(1-0.1498)}{512}\approx 2.49\times 10^{-4},$
$\sigma_{s}=\sqrt{\mathrm{Var}(s)}\approx 0.0158.$
Thus, even though the expected corrupted semantics proportion is about $15\%$, the standard deviation is only around $1.6\%$, confirming that $s$ is tightly concentrated around $p^{k}$ when $N$ is large.
@@ Appendix D A Brief Introduction to SMILES
SMILES (Simplified Molecular Input Line Entry System) (Weininger, 1988) is a notation system used to represent chemical structures in a text format. It encodes molecular information using a series of characters, where atoms are represented by their chemical symbols (e.g., C for carbon, O for oxygen), bonds by symbols like “-” (single bond), “=” (double bond), and “#” (triple bond), and branches are indicated with parentheses. For example, the SMILES representation of ethanol is “CCO”, where “C” stands for carbon atoms and “O” represents an oxygen atom, with a single bond between them. The notation captures the connectivity between atoms, bond types, and sometimes stereochemistry, making it a compact and easily readable way to represent molecular structures. The principle behind SMILES is that it provides a linear representation of the molecule that can be easily interpreted by both humans and computational systems, facilitating tasks like database searching, molecular simulations, and chemical informatics.
@@ Appendix G Training Curves in the Repeated MLM Experiments
Appendix H CUDA-Accelerated Dynamic Programming Framework: Details and Efficiency Analysis
Naively summing over all paths $\mathbf{A}\in\Gamma$ can be prohibitively expensive, since the number of possible paths grows exponentially. To solve this, we use a dynamic programming (DP) approach, which is also adopted by the training algorithms for connectionist temporal classification (Graves et al., 2006; Saharia et al., 2020) and DA-Transformer (Huang et al., 2022; Shao et al., 2022; Huang et al., 2023, 2024). We further adopted a CUDA-accelerated dynamic programming algorithm from DA-Transformer (Huang et al., 2022) to efficiently perform state alignment during pre-training. The core of the algorithm utilizes CUDA (Compute Unified Device Architecture) to parallelize the computations involved in dynamic programming for sequence modeling. The primary objective is to efficiently calculate forward ($\alpha$) and backward ($\beta$) probabilities, which are crucial for evaluating sequence likelihoods and performing gradient-based optimization during model training. By leveraging the parallel processing capabilities of GPUs, the algorithm distributes computations across numerous threads and blocks, significantly reducing runtime compared to traditional CPU-based implementations.
H.1 Dynamic Programming Framework: Forward Probability ($\alpha$) and Backward Probability ($\beta$)
@@ H.2 Gradient Calculations
H.3 Time Complexity Analysis
@@
H.4 Space Complexity Analysis
@@
Appendix I Hyper-Parameter Configuration for Molecular Pre-training
@@
Appendix J Hyper-Parameter Configuration for Molecular Fine-tuning
@@
Appendix K Details of Molecular Fine-tuning Datasets and Baselines
@@
Appendix L Details of the GLUE Benchmark
Below are detailed descriptions of all the GLUE benchmark tasks, which collectively evaluate various aspects of natural language understanding such as entailment, paraphrase detection, sentiment analysis, and grammaticality judgment:
@@
Appendix M Hyper-Parameter Configuration for Textual Pre-training
@@
Appendix N Hyper-Parameter Configuration for Textual Fine-tuning
@@
Appendix O Detailed Case Study
@@
Appendix P More Analytical Experiments of ExLM
P.1 Entropy Analysis on ExLM
@@
P.2 Training Efficiency of ExLM
@@
P.3 Performance of Sparse ExLM